OpenStudy (anonymous):

determine whether the given points is a solution of the equation 1) y=2x+5; (1,7) and 2) 2x+y=2; (-1,4)

OpenStudy (anonymous):

All you have to do is plug the given points into the equation. 1. y = 2x + 5 7 = 2(1) + 5 7 = 2 + 5 7 = 7 True. 2. 2x + y = 2 2(-1) + 4 = 2 -2 + 4 = 2 2 = 2 True. So, the given points both satisfy the equations.

OpenStudy (anonymous):

ohhh I see thank you!:)

OpenStudy (anonymous):

You're welcome! (: